Find the x and y components of r=2a-b if a has components ax=16.6 ay=12.9 where b has components bx=21.5 and by=6.64


1
Expert's answer
2020-09-03T14:10:36-0400

Let the vectors "\mathbf{a}" and "\mathbf{b}" be the following:


"\mathbf{a} = (16.6, 12.9)\\
\mathbf{b} = (21.5, 6.64)"

Then, vector "\mathbf{r}" will be:


"\mathbf{r} = 2\mathbf{a}-\mathbf{b} = 2\cdot(16.6, 12.9) - (21.5, 6.64) = (33.2, 25.8)- (21.5, 6.64)=\\
=(11.7, 19.16)"

Answer. "r_x = 11.7", "r_y = 19.16".